# Accelerating Quantum Monte Carlo Calculations with Set-Equivariant Architectures and Transfer Learning

### v1

@@ 1 Introduction
The explosion of machine learning (ML) in the last two decades has had an extensive and deep impact on condensed matter physics and materials science, with every new advance in architectures and training quickly finding applications in those disciplines [1, 2]. A good illustration is provided by the introduction of advanced regression models as interatomic potentials with ab-initio quality, starting with basic multilayer perceptrons [3] and moving towards sophisticated equivariant graph neural networks [4]. Classification models have been equally prolific, with abundant examples of applications for the analysis of spectroscopic data, as a close analogue of image classification tasks [5]. Likewise, generative models have found many natural uses, some closely related to more mainstream tasks – like the generation of candidate crystalline structures [6], analogous to image or video generation – and some field-specific – such the creation of Boltzmann generators based on normalizing flows to efficiently sample the phase space of classical systems [7].
In the last years, ML strategies have also been deployed to assist in the QMC-based exploration of the rich physics of systems beyond the reach of analytical solutions. Specifically, very general functional forms for regression and methods to train them efficiently have appeared on the scene, which can be used as flexible variational ansätze for a search of the ground state through parametric minimization. The process is similar to – and can therefore reuse part of the machinery developed for – the stochastic minimization of a loss function when training a typical ML regression model, albeit with some important modifications. Firstly, the geometry of the quantum ground state problem favors stochastic reconfiguration [8, 9], a minimization method harnessing more information about the target function than standard ML minimizers like ADAM [10]. More importantly, the quantity being minimized is not a metric of the difference with the ground truth for a predefined set of inputs but a physical quantity. Hence, the concerns about data availability that might be present in an ML regression setting are replaced by the complexity of sampling from the ground-state probability density function and evaluating the expected value of the Hamiltonian and other operators based on those samples. That complexity remains one of the causes of bottlenecks in variational QMC.
@@
In this paper we present an accelerated approach to the calculation of expectation values by making more efficient use of the samples drawn from $\left|\psi\right|^{2}$. The core of our strategy is an architecture developed as a universal, efficient approximation to set-equivariant functions, the set transformer [13], used as the backbone of permutation-invariant regression models. We also show how to build a classifier to detect phase transitions. The calculation of each expectation value using the training architecture is three to four orders of magnitude faster than the direct approach. In addition to analyzing the performance of these models, we discuss how to build them more efficiently by amortizing the training cost over ranges of Hamiltonian parameters, numbers of samples and system sizes: for instance, after pretraining for a chain of $50$ spins, extending the model to $100$ and $150$ spins only requires an additional $70\%$ computational effort.
@@ 2.1 The Set Transformer Architecture
Our goal is to build regression models $f_{\mathcal{O}}\left(\mathcal{S}\subset\mathcal{C}\right)$ to efficiently evaluate MC expectation values of the form expressed by (1) over a random sample $\mathcal{S}$ from the computational basis $\mathcal{C}$. We also aim to develop classification models taking $\mathcal{S}$ as their input and emitting a verdict about the phase of the system (i.e., to answer whether it is ferromagnetic or not). A key design consideration of such models is that they treat $\mathcal{S}$ as a set, i.e., the answer can only depend on which elements of $\mathcal{C}$ are included in $\mathcal{S}$. Therefore, the functional form of the regressions and classifications must be completely permutation invariant and allow for a variable number of samples in $\mathcal{S}$.
@@
A more practically oriented approach to permutation invariance that still results in an universal approximator of set-invariant functions is the set transformer [13], derived from the extremely successful transformer architecture [17]. The set transformer can be described starting from a conventional multihead dot-product attention block with LayerNorm normalization [18] but without any positional encoding [19] or dropout [20]. We will denote that block as $\operatorname{MAB}\left(\bm{X},\bm{Y}\right)$, where $\bm{X}$ and $\bm{Y}$ are two arrays of $n$ $d$-dimensional tokens each, with $\bm{X}$ being used to generate the queries and $\bm{Y}$ being used to generate the keys and values of the attention mechanism. This basic $\operatorname{MAB}$ block is then used to build two different kinds of higher-level components: permutation-equivariant set-attention blocks and permutation-invariant pooling blocks. A full set transformer contains an encoder pipeline in the form of a stack of set-equivariant blocks followed by a set-invariant decoder stage producing the final output.
A set-equivariant block can be straightforwardly built by turning the multihead attention block into a self-attention block by short-circuiting its inputs $\operatorname{SAB}\left(\bm{X}\right)=\operatorname{MAB}\left(\bm{X},\bm{X}\right)$; in fact, it is this well-known property of the dot-product self attention that motivates the introduction of positional encodings in applications where neighborhood relations between input tokens are important and therefore set equivariance is undesirable. However, the scalability of this simple $\operatorname{SAB}$ is poor and a more performance-oriented alternative is the induced set attention block $\operatorname{ISAB}_{m}\left(\bm{X}\right)=\operatorname{MAB}\left[X,\operatorname{MAB}\left(\bm{I},\bm{X}\right)\right]$, where $\bm{I}$ is a set of $m<n$ trainable inducing points.
@@
The building blocks of the set transformer are schematically represented in Figure 1; following the original recipe of the set transformer [13], in this work we employ implementations of the set transformer where each processing component is simply a non-linear unit, more specifically a ReLU [21].
@@
Figure 1: Diagrams illustrating the structure of the main blocks of the set transformer. As described in the text, $\bm{S}$ and $\bm{I}$ are also trainable inputs.
@@ 2.2 Our Regression and Classification Models
We use two specific set transformer models, one for regression and one for classification. Our classifier contains an encoder stage made up of a $5$-head $\operatorname{SAB}$ followed by a $10$-head SAB, a decoder stage comprising a $5$-head, single-seed $\operatorname{PMA}$ followed by two $10$-head $\operatorname{SAB}$s with a LayerNorm between them, a linear layer to generate the logits, and a SoftMax layer to generate the final output. When we apply this model to label phases as antiferromagnetic, ferromagnetic or paramagnetic, we obtain better results by using a sequence of two binary classifiers than by resorting to a direct ternary classifier. The first of those classifiers sets aside paramagnetic states; the remaining ones are passed on to the second classifier, which then discriminates between ferromagnetic and antiferromagnetic states.
@@ 2.3 Hamiltonian, Ansatz and Ground-State Search
To find the ground state of the Hamiltonian in (2) we optimize the parameters of a neural-quantum-state ansatz. In particular, we choose a flexible functional form based on the vision transformer (ViT [28]), another widely used specialization of the transformer architecture. In our model, the projection of state $\left\lvert\psi\right\rangle$, represented by the ansatz, on element $\lvert s\rangle$ of the computational basis is determined through the following steps:
@@ 2.4 Data Generation and Training Procedures
To perform the training of the classifier architecture, we use only states with $\alpha\in\left\{0,3.2,6\right\}$. We augment the training set, increasing the number of its elements by a factor of five, through a subsampling aggregation scheme where in each iteration we choose a member of the original training set, draw a uniform random integer between $1$ and $1024$ and then sample a random subset with that cardinal. This same technique is applied to the validation and test sets. The learning process is driven by an Adam optimizer with a fixed value of $5\times 10^{-5}$ for the learning rate and had a total duration of $500$ epochs.
@@
$\bm{\hat{\bm{m_{z}}}}=\begin{cases}\frac{1}{N}\sum_{i=1}^{N}\sigma_{z}^{i}&\quad J<0,\\ \frac{1}{N}\sum_{i=1}^{N}\left(-1\right)^{i}\sigma_{z}^{i}&\quad J>0.\end{cases}$
@@ 2.5 Transfer Learning
To make that kind of workflow possible, we need the ML model $f_{\mathcal{O}}\left(\mathcal{S}\subset\mathcal{C}\right)$ to accept sets $\mathcal{S}$ with samples of different length (although each $\mathcal{S}$ is still homogeneous). Note that this is a different and additional requirement than just being able to accept sets with different cardinals, which is accomplished using attention masks. To allow the model to work with different numbers of spins we allow three spin values in the input, $\left\{0,1,2\right\}$, with 0 and $1$ playing the roles previously assigned to $\pm 1$ and $2$ representing missing spins. We then add a trainable embedding layer after the input to map those to more optimal values for the ML models. We do not use masking along this dimension, but let the networks learn how to interpret the ternary input instead. Specifically, we start training the network with the data from states with 50 spins. Once this step is accomplished, we freeze the gradients of the main block of the architecture, and train only the final heads of the network to compute states composed of 100 and 150 spins. Both of these subsequent training steps take about $1/3$ less time than the first one.
@@ 3 Results
Closely related to this classification is the regression model for $\left<m^{2}\right$>, whose results are shown in the left panel of Figure 3. The transitions among the three phases are again immediately visible and agree with previous results. Furthermore, the logarithm of the difference between the predicted results and those calculated with QMC is shown in the right side of Figure 3. The model seems to perform worst in the ferromagnetic phase, where errors can reach $\sim 15\%$, than in the other two phases of the system, and are primary located within certain bands, e.g. $\alpha\in\{1,2,6\}$. This correlates with the worse quality of the training data in the ferromagnetic region.
@@
As explained in the previous section, for the second Rényi entanglement entropy, the set transformer was trained on a set of states where each pair $\{\alpha,J/h\}$ has five realizations, but each one of them is labeled by their mean entanglement entropy, instead of the individual values. This is done with the intention of averaging out the numerical fluctuations that arise from the QMC calculations, and at the same time enables us to get averaged outputs of a single state used as input once the regressor is trained, which drastically reduces the computation time. Furthermore, we employ a transfer-learning technique, as explained in Section 2.5, in order to reduce the training cost for the larger system sizes once the model has been trained on the shortest chain. The left panel of Figure 5 shows the predictions of the set transformer for the states of $\alpha=2.5$ for system sizes of $N=50,100,150$, where we also used a Savitzky-Golay filter [39], which allows us to better capture the trend with increased size as we approach the critical point. To further understand the increase of the entanglement with the system size at the phase transition, we study their relationship. In the right panel of Figure 5 we show the predicted logarithmic volume law calculated with the results of the set transformer. This result can be compared with those found in the literature, such as in [26], where the volume law reported was $\langle S_{2}\rangle=0.19\log(N)+0.01$.

### v2

@@ 1 Introduction
The explosion of machine learning (ML) in the last two decades has had an extensive and deep impact on condensed matter physics and materials science, with every new advance in architectures and training quickly finding applications in those disciplines [1, 2]. A good illustration is provided by the introduction of advanced regression models as interatomic potentials with ab-initio quality, starting with basic multilayer perceptrons [3] and moving towards sophisticated equivariant graph neural networks [4]. Classification models have been equally prolific, with abundant examples of applications for the analysis of spectroscopic data, as a close analogue of image classification tasks [5]. Likewise, generative models have found many natural uses, some closely related to more mainstream tasks —like the generation of candidate crystalline structures [6], analogous to image or video generation— and some field-specific —such the creation of Boltzmann generators based on normalizing flows to efficiently sample the phase space of classical systems [7].
In the last years, ML strategies have also been deployed to assist in the QMC-based exploration of the rich physics of systems beyond the reach of analytical solutions. For instance, classifiers have been employed to detect phase transitions in problematic many-fermion systems [8, 9] and in a more general setting [10]. Moreover, very general functional forms for regression and methods to train them efficiently have appeared on the scene [11], which can be used as flexible variational ansätze for a search of the ground state through parametric minimization. The process is similar to —and can therefore reuse part of the machinery developed for— the stochastic minimization of a loss function when training a typical ML regression model, albeit with some important modifications. Firstly, the geometry of the quantum ground state problem favors stochastic reconfiguration [12, 13], a minimization method harnessing more information about the target function than standard ML minimizers like ADAM [14]. More importantly, the quantity being minimized is not a metric of the difference with the ground truth for a predefined set of inputs but a physical quantity. Hence, the concerns about data availability that might be present in an ML regression setting are replaced by the complexity of sampling from the ground-state probability density function and evaluating the expected value of the Hamiltonian and other operators based on those samples. That complexity remains one of the causes of bottlenecks in variational QMC.
@@
In this paper we present an accelerated approach to the calculation of expectation values by making more efficient use of the samples drawn from $\left|\psi\right|^{2}$. A critical feature of such approaches is that the estimate of an expectation value cannot depend on the order in which the $\left\lvert\nu\right\rangle$ are drawn from $\left|\psi\right|^{2}$ or fed to the model, i.e., it must be permutation invariant. Moreover, they must be able to work on collections of $\left\lvert\nu\right\rangle$ with different cardinality. Those are the characteristics of a set-invariant architecture. We have therefore based our strategy on an architecture developed as a universal, efficient approximation to set-equivariant functions, the set transformer [17]. We use it as the backbone of permutation-invariant regression models. We also show how to build a classifier based on the set transformer to detect phase transitions. The calculation of each expectation value using the training architecture is three to four orders of magnitude faster than the direct approach. In addition to analyzing the performance of these models, we discuss how to build them more efficiently by amortizing the training cost over ranges of Hamiltonian parameters, numbers of samples and system sizes: for instance, after pretraining for a chain of $50$ spins, extending the model to $100$ and $150$ spins only requires an additional $70\%$ computational effort.
@@ 2.1 The Set Transformer Architecture
Our goal is to build regression models $f_{\mathcal{O}}\left(\mathcal{S}\subset\mathcal{C}\right)$ to efficiently evaluate MC expectation values of the form expressed by (1) over a random sample $\mathcal{S}$ from the computational basis $\mathcal{C}$. We also aim to develop classification models taking $\mathcal{S}$ as their input and emitting a verdict about the phase of the system (i.e., to answer whether it is ferromagnetic or not). A key design consideration of such models is that they treat $\mathcal{S}$ as a set, i.e., the answer can only depend on which elements of $\mathcal{C}$ are included in $\mathcal{S}$. Therefore, the functional form of the regressions and classifications must allow for a variable number of samples in $\mathcal{S}$ and be completely permutation invariant, i.e., $f_{\mathcal{O}}\left[\tau\left(\mathcal{S}\right)\right]=f_{\mathcal{O}}\left(\mathcal{S}\right)$ for any permutation $\tau\in\operatorname{Sym}\left(\mathcal{S}\right)$.
@@
A more practically oriented approach to permutation invariance that still results in an universal approximator of set-invariant functions is the set transformer [17], derived from the extremely successful transformer architecture [21]. The set transformer can be described starting from a conventional multihead dot-product attention block with LayerNorm normalization [22] but without any positional encoding [23] or dropout [24]. We will denote that block as $\operatorname{MAB}\left(\bm{X},\bm{Y}\right)$, where $\bm{X}$ and $\bm{Y}$ are two arrays of $n$ $d$-dimensional tokens each, with $\bm{X}$ being used to generate the queries and $\bm{Y}$ being used to generate the keys and values of the attention mechanism. This basic $\operatorname{MAB}$ block is then used to build two different kinds of higher-level components: set-attention blocks ($\operatorname{SAB}$) and pooling-by-multihead-attention ($\operatorname{PMA}_{k}$) blocks. While $\operatorname{PMA}_{k}$ blocks are set invariant, $\operatorname{SAB}$s are set equivariant, i.e., $\operatorname{SAB}\left[\tau\left(\bm{X}\right)\right]=\tau\left[\operatorname{SAB}\left[\bm{X}\right]\right]\;\forall\tau\in\operatorname{Sym}\left(\bm{X}\right)$. A full set transformer contains an encoder pipeline in the form of a stack of set-equivariant blocks followed by a decoder stage producing the final output. In our models we design the decoder stage to be set invariant, thus making the whole transformer invariant as well.
A set-equivariant block can be straightforwardly built by turning the multihead attention block into a self-attention block by short-circuiting its inputs $\operatorname{SAB}\left(\bm{X}\right)=\operatorname{MAB}\left(\bm{X},\bm{X}\right)$; in fact, it is this well-known property of the dot-product self attention that motivates the introduction of positional encodings in applications where neighborhood relations between input tokens are important and therefore set equivariance is undesirable. However, the scalability of this simple $\operatorname{SAB}$ is limited and a more performance-oriented alternative is the induced set attention block $\operatorname{ISAB}_{m}\left(\bm{X}\right)=\operatorname{MAB}\left[X,\operatorname{MAB}\left(\bm{I},\bm{X}\right)\right]$, where $\bm{I}$ is a set of $m<n$ trainable inducing points.
@@
For the system sizes employed in this work, we do not need to resort to the ISAB. Our set transformers are built entirely our of self-attention and pooling-by-self-attention blocks. The whole set of building blocks is schematically represented in Figure 1; following the original recipe [17], in this work we employ implementations of the set transformer where each processing component is simply a non-linear unit, more specifically a ReLU [25].
@@
Figure 1: Structure of the main building blocks of the Set Transformer. The Multihead Attention Block (MAB) is the core component, defined as a standard multihead dot-product attention with LayerNorm and ReLU, but without positional encoding or dropout. The Induced Set Attention Block (ISAB) is an alternative to self-attention that preserves the set equivariance but improves scaling; it is constructed as $\operatorname{ISAB}_{m}\left(\bm{X}\right)=\operatorname{MAB}\left[\bm{X},\operatorname{MAB}\left(\bm{I},\bm{X}\right)\right]$ with $m<n$ trainable inducing points $\bm{I}$. The Pooling by Multihead Attention (PMA) block defines the set-invariant decoder stage, specified as $\operatorname{PMA}_{k}\left(\bm{X}\right)=\operatorname{MAB}\left[\bm{S},\operatorname{FF}\left(\bm{X}\right)\right]$, where $\bm{S}$ are $k$ trainable seed vectors and $FF$ is a feed-forward block.
@@
A set-transformer-based approach is model-agnostic, and can be applied in other kinds of QMC. The only requirement is access to configurations drawn from a sampling measure, be they fermionic samples, occupation snapshots, Hubbard–Stratonovich slices, or equal-time Green matrices. For instance, in Ref. [8], Broecker et al. demonstrated that a CNN trained on equal-time Green matrices produced by auxiliary-field QMC can locate fermionic phase transitions even when conventional sign-reweighted estimators fail. Their key idea is to “state function” directly from QMC configurations, thereby bypassing explicit sign reweighting in the final prediction. This complements our set-equivariant approach: a set transformer can be trained with exactly the same sampled objects and perform phase classification without computing observables, thus inheriting the same practical insensitivity to the sign problem for classification tasks.
@@ 2.2 Our Regression and Classification Models
We use two specific set transformer models, one for regression and one for classification. Our classifier contains an encoder stage made up of a $5$-head $\operatorname{SAB}$ followed by a $10$-head SAB, a decoder stage comprising a $5$-head, single-seed $\operatorname{PMA}$ followed by two $10$-head $\operatorname{SAB}$s with a LayerNorm between them, a linear layer to generate the logits, and a SoftMax layer to generate the final output. We do not include any $\operatorname{ISAB}$ block since the quadratic scaling of attention does not become a problem for these system sizes. When we apply this model to label phases as antiferromagnetic, ferromagnetic or paramagnetic, we obtain better results by using a sequence of two binary classifiers than by resorting to a direct ternary classifier. The first of those classifiers sets aside paramagnetic states; the remaining ones are passed on to the second classifier, which then discriminates between ferromagnetic and antiferromagnetic states.
@@ 2.3 Hamiltonian, Ansatz and Ground-State Search
To find the ground state of the Hamiltonian in (2) we optimize the parameters of a neural-quantum-state ansatz. In particular, we choose a flexible functional form based on the vision transformer (ViT [32, 11]), another widely used specialization of the transformer architecture. In our model, the projection of state $\left\lvert\psi\right\rangle$, represented by the ansatz, on element $\lvert s\rangle$ of the computational basis is determined through the following steps:
@@ 2.4 Data Generation and Training Procedures
To perform the training of the classifier architecture, we use only states with $\alpha\in\left\{0,3.2,6\right\}$. We augment the original training set, $T$, increasing the number of its elements by a factor of five, through the following subsampling aggregation scheme:
  1. 1.
    At each iteration, we draw a random element $t\in T$, and a uniform random integer $c\in\left[1,1024\right]$.
  2. 2.
    We extract a random subsample of $c$ elements from the input part of $t$.
  3. 3.
    We add that subsample, along with the output part of $t$, to the new training set.
This procedure not only enlarges, but also enriches the set with elements of different cardinality, whose correct treatment is one of the key features we are trying to promote in our models. This same technique is applied to the validation and test sets. The learning process is driven by an Adam optimizer with a fixed value of $5\times 10^{-5}$ for the learning rate and had a total duration of $500$ epochs.
@@
$\bm{\hat{\bm{m_{z}}}}=\begin{cases}\frac{1}{N}\sum_{i=1}^{N}\sigma_{z,i}&\quad J<0,\\ \frac{1}{N}\sum_{i=1}^{N}\left(-1\right)^{i}\sigma_{z,i}&\quad J>0.\end{cases}$
@@
Figure 2: Workflow of the transfer learning scheme. The model is first trained on data from the smaller system. A new model then inherits a subset of the trained parameters, which are frozen to prevent further updates. The remaining parameters are subsequently trained on data from the larger system. “Dense” refers to a fully connected block, whose precise form will be different for regressions and classifications.
Figure 3: Accuracy of the predictions with and without transfer learning for $N=100$ and $\alpha=2.5$. The standard error over the validation set is computed as $\delta m^{2}_{z}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(m_{z,i}^{2}-\tilde{m}_{z,i}^{2}\right)^{2}}$ where $n$ is the number of data points in that set, $m_{z,i}^{2}$ are the fluctuations of the magnetization calculated via QMC with original variational states and $2048$ samples, and $\tilde{m}_{z,i}^{2}$ are the set-transformer predictions for the same quantities.
@@ 2.5 Transfer Learning
To make that kind of workflow possible, we need the ML model $f_{\mathcal{O}}\left(\mathcal{S}\subset\mathcal{C}\right)$ to accept sets $\mathcal{S}$ with samples of different length (although each $\mathcal{S}$ is still homogeneous). Note that this is a different and additional requirement than just being able to accept sets with different cardinals, which is accomplished using attention masks. To allow the model to work with different numbers of spins we allow three spin values in the input, $\left\{0,1,2\right\}$, with 0 and $1$ playing the roles previously assigned to $\pm 1$ and $2$ representing missing spins. We then add a trainable embedding layer after the input to map those to more optimal values for the ML models. We do not use masking along this dimension, but let the networks learn how to interpret the ternary input instead.
Figure 2 illustrates the transfer learning workflow employed in this study. The network is first trained on data from the smaller system (50 spins). After this initial stage, the gradients of the main architectural block are frozen to prevent further updates, and only the final heads of the network are retrained from scratch using data from the larger systems (100 and 150 spins). Each of these subsequent training stages requires approximately one third less time than the initial training.
Although transfer learning does not enable the exploration of entirely new system sizes, as data from the target systems is still required, it improves efficiency by reducing the amount of data needed compared to training a network from scratch. Figure  3 illustrates this fact through a direct comparison of the accuracies of networks trained to predict the fluctuations in the magnetization, with or without transfer learning, as a function of the fraction of the available data used in their training.
@@ 3 Results
Closely related to this classification is the regression model for $\left<m^{2}\right$>, whose results are shown in the left panel of Figure 5. The transitions among the three phases are again immediately visible and agree with previous results. Furthermore, the logarithm of the difference between the predicted results and those calculated with QMC is shown in the right side of Figure 5. The model seems to perform worst in the ferromagnetic phase, where errors can reach $\sim 15\%$, than in the other two phases of the system, and are primary located within certain bands, e.g. $\alpha\in\{1,2,6\}$. This correlates with the worse quality of the training data in the ferromagnetic region. The training process takes approximately 15 minutes for the first system size and about 11 minutes for each subsequent one. Once training is complete, the evaluation of an individual expectation value requires only $0.04\text{\,}\mathrm{s}$, which amounts to a speed-up of roughly $\times 100$ for $N=50$ and $\times 1800$ for $N=100$.
@@
As explained in the previous section, for the second Rényi entanglement entropy, the set transformer was trained on a set of states where each pair $\{\alpha,J/h\}$ has five realizations, but each one of them is labeled by their mean entanglement entropy, instead of the individual values. This is done with the intention of averaging out the numerical fluctuations that arise from the QMC calculations, and at the same time enables us to get averaged outputs of a single state used as input once the regressor is trained. Furthermore, we employ a transfer-learning technique, as explained in Section 2.5, in order to reduce the training cost for the larger system sizes once the model has been trained on the shortest chain. The left panel of Figure 7 shows the predictions of the set transformer for the states of $\alpha=2.5$ for system sizes of $N=50,100,150$, where we also used a Savitzky-Golay filter [43], which allows us to better capture the trend with increased size as we approach the critical point. To further understand the increase of the entanglement with the system size at the phase transition, we study their relationship. In the right panel of Figure 7 we show the predicted logarithmic volume law calculated with the results of the set transformer. This result can be compared with those found in the literature, such as in [30], where the volume law reported was $\langle S_{2}\rangle=0.19\log(N)+0.01$.